Use the binomial theorem to expand the binomial...
(3x^4-2)^4

Answer by Nate(3500)

(A + B)^4 = 1*A^4 + 4*A^3*B^1 + 6*A^2*B^2 + 4*A^1*B^3 + 1*B^4
(3x^4-2)^4 = ((3x^4) + (-2))^4 ... A = 3x^4 and B = -2
Now, substitute and simplify for your answer.